OpenStudy (anonymous):

I am trying to figure out what the difference between 2 heights are. Ann is 67 5/6 inches tall and Mary is 71 1/7 inches tall. How much taller is marry than ann. i have tried 2.8367 but that was wrong

OpenStudy (anonymous):

(71*7 + 1 )/7 - (67*6 +5) / 6) = 139/42 inches taller.

OpenStudy (anonymous):

Is 4 inches and 29/49 correct?

OpenStudy (anonymous):

\[\left(71\frac{1}{7}\right)-\left(67\frac{5}{6}\right)=3\frac{13}{42}=3.30952 \]